ID Let f(x) = /2x=2 | (a) Find f(x) as a piecewise function (6) Graph y = f(x) (c) state the domain of f and the domain oftanx 7 Let y= (a) Find (6) Find the equation of the tangent line at (74 y Elf 8 X3 Prove lim (5-) = 4 (a) write the e-s defi
Answer #1

(11) Let L = lim tum 4x Cas 3x Cå forom) X70 Sin 52 so we care apply L’Hospital rule. L = lim 4 sec 4x. Cesgx + 3tan4x (-Sin13 we have 1 (GB 3+10x) -3 -0% y 3/ **+10% put x3 + 10x So from ) cu -> (14) 9 M =zē 3 dy are 1 ze ze of Again from du = (3x+1 16 [ = s (7² 2 T (x - 1 ay π T 7. A x(JZ) - 1 or, y - 17 2 T2 16(1-2) 272 8(7-2) 72 o the eqreation of the tangent line